# AURA-CVC: Autonomous Ultrasound-guided Robotic Assistance for Central Venous Catheterization

### v1

@@ 1 Introduction
Central venous catheterization (CVC) is a critical procedure used in intensive care, trauma resuscitation, major surgeries, parenteral nutrition delivery, and hemodialysis. Over 7 million CVCs are placed annually in the U.S. alone [1]. It is typically performed at one of three primary sites: the internal jugular (IJ), subclavian, or femoral veins, with IJ as a preferred site due to its superficial location, consistent anatomy, and lower complication risk. Moreover, ultrasound-guided CVC has become clinical standard due to its improved safety [2]. The procedure begins with scanning the target region via ultrasound to identify vessels, followed by estimating the insertion point and needle trajectory. Major complications are observed in approximately 3% of 1000 CVC placements, with rates reported as $\sim 4.4$ for pneumothorax, $\sim 16.2$ for arterial puncture, and $\sim 2.8$ for arterial cannulation [3]. Although ultrasound guidance reduces complications, maintaining continuous visualization of both the vessel and advancing needle in the narrow imaging plane remains a major challenge [4]. Even experienced practitioners have to carefully balance probe positioning and needle advancement. These challenges highlight the need for advanced technologies to reduce workload and improve safety.
@@
Robot-assisted ultrasound (RUS) has emerged as a promising solution to enhance precision, minimize human error, and improve procedural outcomes in various diagnostic and interventional applications [5, 6, 7]. Several handheld and semi-automated needle insertion devices have been developed to optimize workspace, accuracy, and force resistance [8, 9, 10]. Among them, CathBot [10] allows clinicians to initiate insertion by pushing a handle, with the device completing cannulation autonomously. While this was effective for peripheral vascular access procedures on forearm, it has limitations in CVC due to anatomical complexities at neck, chest, or groin insertion sites. Later, the focus shifted towards enhancing the dexterity of these mechanisms using advanced robotics, imaging, and sensor guidance for needles. Hong et al. [11] proposed a percutaneous needle insertion robot that can compensate for involuntary patient motion using real-time visual-servoing. Neubach et al. [12] presented a closed-loop robotic controller for flexible needle steering by calculating the needle tip position from ultrasound images, while accounting for tissue displacement from a stiffness model. Xu et al. [13] proposed a slice guidance method, which interpolated preoperative scans in real-time to compensate for the motion of liver during needle placement. Chatelain et al. [14] used particle filtering to track the flexible needle for real-time closed-loop control during targeting. While these approaches [11, 12, 13, 14] demonstrate promising closed-loop control and organ motion compensation techniques, they lack integrated intra-operative planning for a fully autonomous procedure.
@@
Zettinig et al. [15] annotated pre-interventional plans in CT/MRI and proposed a 3D visual servoing framework with multi-modal registration to dynamically update the needle target with anatomical motion. Kojcev et al. [16] developed a dual-arm ultrasound-guided system integrating force-controlled image acquisition, registration, vision-based control, and needle tracking. However, intra-operative planning still relied on manual region selection. Bell et al. [17] and Gubbi et al. [18] demonstrated photoacoustic-based (rather than ultrasound) visual servoing of neeedle. Cheng et al. [19] used electrical impedance sensing to detect vascular entry. In these cases, no intra-operative planning with CT or MRI registration was used. Koskinopoulou et al. [20] combined ultrasound and bioimpedance in a dual-arm system for robust vessel localization and venipuncture detection. While the above-discussed systems [15, 17, 18, 16, 19, 20] successfully demonstrated end-to-end procedures, scanning was either initialized from a predefined pose or relied on medical experts to manually review intra-operative CT/MRI images.
@@
    An end-to-end robotic pipeline to perform key CVC steps: intra-operative planning, scanning, anatomical localization, and needle placement.
@@
    A method to predict the optimal scanning region and trajectory using only a depth image from an RGB-D camera.
@@
    Automated robotic scanning and vascular reconstruction for localization of vein and artery, followed by supervised needle guidance.
@@
    Validation on a near-realistic CVC phantom for IJ vein access across 10 diverse clinical scenarios.
@@ 2.1 Experimental testbed
The experimental testbed (Fig. 1) includes: (1) a $6-$DoF UR10e robotic arm (Universal Robots, Denmark) with $10$ kg payload and inbuilt force-torque sensor, (2) a C3 HD3 curved ultrasound probe (Clarius, Canada) mounted via an extended attachment for real-time imaging. Although linear probes are preferred for CVC, we used a convex probe for its wider field of view and availability during prototyping. (3) An Ubuntu PC and a laptop, with NVIDIA RTX 4090 (16GB) and 1090 (12GB) for system control and deep learning inferences, respectively, (4) an Intel RealSense D415 RGB-D camera (California, USA) for spatial localization, (5) a custom 2-DoF needle insertion mechanism, adapted from the design in [21], providing linear and angular control, compatible with any robotic arm with adequate payload capacity, and equipped with an 18-gauge, 15 cm percutaneous thin-wall bevel-tipped needle (Cook Medical, USA), (6) The CVC Blue Phantom (CAE, USA) for validation, offering highly realistic upper thorax and neck anatomy, including major vessels and anatomical landmarks commonly used in CVC. Note that the communication across all components of the setup is established via ROS2 (Robot Operating System).
Let $\boldsymbol{T}_{1}^{2}$ denote the homogeneous transformation matrix that expresses the pose of frame 1 relative to frame 2. The frames used in the system are: robot’s base frame (B), end-effector (E), camera (C), probe tip (P), remote center-of-motion (R), and ultrasound image (U) (see Fig. 1). The 3D position in world coordinate system ($p_{B}$) of a given pixel in the 2D ultrasound image ($p_{I}$) is calculated using:
@@
where $\boldsymbol{T}_{U}^{I}$ is the transformation from the ultrasound image origin (top left corner) to the pixel values in the image and is given by
$\boldsymbol{T}_{U}^{I}=\begin{bmatrix}S_{x}&0&0&0\\ 0&S_{y}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\begin{bmatrix}I_{x}\\ I_{y}\\ 0\\ 1\end{bmatrix}$ (2)
where $S_{x}=S_{y}=0.15436~{}mm$ is the pixel spacing (scaling factor) provided by the Clarius Mobile Health software development kit and $(I_{x},I_{y})$ is the given pixel value in $x-$ and $y-$direction of the image.
@@ 2.2 Pipeline
Our pipeline is shown in Fig. 2. During initialization, system calibration was performed by estimating $\boldsymbol{T}_{C}^{B}$ via hand-eye calibration and $\boldsymbol{T}_{U}^{E}$ using the classical $AX=XB$ method [22], with a cross-wire phantom in a water bath. The IJ vein and carotid artery region was then autonomously identified from depth images, followed by robotic ultrasound scanning, vessel segmentation, 3D reconstruction, insertion point selection, and closed-loop needle guidance for puncture.
@@ 2.3 Intra-operative planning
Clinicians often define a triangular region using the suprasternal notch (SN), sternocleidomastoid (SCM) midpoint, and lateral clavicle (LC), known as Sedillot’s triangle, to guide IJ vein access [23]. Inspired by this, we used SN, LC, and anterior hyoid bone tip (HT) to define a similar region, with HT chosen for its prominence and absence of SCM in the phantom. This region still ensured IJ vein’s presence. To identify these landmarks, we used a Dense-UNet deep learning model, as described below.
@@ 2.3.1 Anatomical landmark prediction network
The anatomical landmark prediction network required a depth image of neck and torso region of the body to predict the landmarks. First, a synchronized depth and color images of the environment was captured by the depth camera. An HSV-based thresholding then segmented the region-of-interest i.e. phantom’s head and torso, followed by depth filtering for a refined 3D point cloud. The segmented RoI point cloud was provided as input to the network, as shown in Fig. 3(a).
Following the approach in [24], we trained a Dense-UNet [25] to predict the location of internal landmarks from computed depth images. The network architecture consisted of an encoder and decoder, each with four dense blocks of four convolutional layers (Fig. 3(b)). Batch normalization and ReLU activation were applied after each convolutional layer, except for the final layer.
@@
The dataset of 4,912 depth images was used to train the Dense-UNet model, consisting of real human subjects with diverse body habitus, sex, and skin tone profiles. Each image was paired with manually annotated 3D anatomical landmarks by trained technicians using co-registered RGB-D data and verified against physical anthropometric measurements to ensure accuracy. The network was trained for 300 epochs, minimizing the Adaloss function using the using the Adam optimizer (learning rate = 0.001, batch size = 32). The network predicted 33 landmarks, using additional ones for regularization to enhance accuracy of the 3 landmarks critical to CVC.
@@ 2.3.2 Scanning region and path estimation
The scanning region was defined by triangular region formed by 3 predicted landmarks (SN, LC, and HT). This region was projected onto the phantom surface using the Open3D Python library, as shown in Fig. 3(c). To scan within this region, we defined straight-line paths from the triangle’s base (SN–LC) to its apex (HT), aligning with the IJ vein’s orientation. Initial paths connected HT to LC and HT to SN. Additional scanning paths were created by dividing the SN–LC line into equal intervals and connecting lines from HT to these points, yielding a dense scan pattern for accurate vessel detection (Fig. 3(d)). Scanning was initiated along the midpoint-to-HT path because this path aligned with the expected anatomical course of the IJ vein, offering a clinically representative initial view. Additionally, it provided a reliable starting point across different patient positioning due to its geometric symmetry.
@@ 2.4 Robot motion planning
Once the scanning region was identified, the pipeline generated a robot trajectory for scanning and localizing the vein-artery, as illustrated in Fig. 4. After robotic scanning, the vein was localized and centered in the ultrasound image to ensure consistent spatial referencing. Then the robot was re-oriented for sagittal plane insertion, as it offered full needle shaft visibility. A collision check was performed using the CAD model and phantom point cloud. If a collision was detected in the sagittal orientation, the robot iteratively adjusted its pivot angle in transverse view about the point of contact by $5^{\circ}$ until a collision-free pose was found. Upon confirming clearance, the robot transitioned to the sagittal plane for needle insertion.
@@ 2.4.1 Scanning trajectory
Each pose transform $\boldsymbol{P}_{*}$ is defined as $[\overrightarrow{\boldsymbol{x}},\overrightarrow{\boldsymbol{y}},\overrightarrow{\boldsymbol{z}},\overrightarrow{\boldsymbol{t}}]$, where $\overrightarrow{\boldsymbol{t}}$ is the translation (or position) vector and $\overrightarrow{\boldsymbol{x}},\overrightarrow{\boldsymbol{y}},\overrightarrow{\boldsymbol{z}}$ are unit vectors for orientation. The start and end of the scanning path define $\boldsymbol{P}_{0}$ and $\boldsymbol{P}_{n}$, respectively. The normal $\overrightarrow{\boldsymbol{z}}$ of phantom was computed using Open3D and aligned with the end-effector’s normal $\overrightarrow{\boldsymbol{z}_{e}}$ using:
@@
$\overrightarrow{\boldsymbol{y}}=\frac{\overrightarrow{\boldsymbol{z}}\times\overrightarrow{\boldsymbol{x}}}{||\overrightarrow{\boldsymbol{z}}\times\overrightarrow{\boldsymbol{x}}||}$ (5)
Intermediate poses were interpolated between $\boldsymbol{P}_{0}$ and $\boldsymbol{P}_{n}$ using the UR driver. ROS2 and URScript executed the trajectory in hybrid force-position mode with $5N$ force along the probe’s $z$-axis. During scanning, a rosbag recorded poses and images pairs $(\boldsymbol{P}_{i},\boldsymbol{I}_{i})$ for each frame, as given in eq. (6), to enable accurate vessel reconstruction.
@@ 2.4.2 Identifying optimal insertion plane
The acquired 2D US frames were segmented to reconstruct vessels and plan needle insertion. A deep learning model from prior work [26] was used to extract vessel center-points and areas from each frame. Ideally, insertion should occur where both vessels (vein and artery) are clearly visible for confident and precise targeting. To determine this, the robot insertion pose ($\boldsymbol{\mathcal{P}}$) was selected from $\boldsymbol{P}_{i},\cdots,\boldsymbol{P}_{n}$ as the one corresponding to the maximum combined area of the two vessels. This selection criterion is illustrated in Fig. 5. The robot was finally moved to pose $\boldsymbol{\mathcal{P}}$, and the corresponding ultrasound frame was re-segmented.
@@ 2.4.3 Vein-artery localization
Veins are difficult to distinguish from arteries, as both appear morphologically similar in ultrasound images. While clinicians typically use Doppler ultrasound to differentiate them based on blood flow, this method requires manual tuning, favorable imaging conditions, and expertise, making it less practical in portable or low-cost devices where Doppler functionality or detectable flow is absent. To address this, our system employs a compression-based method that exploits differences in vascular compliance, whereby veins exhibit greater deformation than arteries under low external pressure.
The robot applied controlled compression by gradually increasing the contact force from $5$ N to $10$ N, then decreasing it back to $5$ N, while recording an ultrasound video. Segmented vessel areas and centroids were extracted from the recorded sequence. Then, KMeans clustering was used, which grouped vessel centroids into two clusters based on their spatial proximity. Area changes across the compression cycle were computed for each cluster, with the one showing greater deformation identified as the vein, and the other as the artery. The centroid of the vein cluster was stored as a reference. Later, at the insertion pose, the robot captured a new ultrasound frame, segmented the vessels, and identified the vein by finding the centroid closest to the saved reference, with the remaining vessel labeled as the artery.

### v2

@@ 1 Introduction
Central venous catheterization (CVC) is widely performed in intensive care, trauma, major surgery, parenteral nutrition, and hemodialysis, with over 7 million annual placements in the U.S. [kehagias2023central]. It is typically performed at the internal jugular (IJ), subclavian, or femoral veins, with IJ preferred for its superficial location, consistent anatomy, and reduced risk. Ultrasound-guided CVC is now the clinical standard [govindan2018peripherally]. The procedure involves ultrasound scanning to identify vessels, estimating insertion point, and advancing the needle. Major complications occur in 3% of placements (i.e., $4.4$ pneumothorax, $16.2$ arterial puncture, $2.8$ arterial cannulation per 1000 cases) [teja2024complication]. Despite ultrasound guidance, continuous simultaneous visualization of a vessel and needle is difficult due to narrow imaging planes and vessel “roll-away” during advancement. These challenges highlight the need for advanced technologies to reduce workload and improve safety.
@@
Robot-assisted ultrasound (RUS) has emerged as a promising solution to enhance precision, minimize human error, and improve outcomes in diagnostic and interventional applications [biswas2023recent, raina2021comprehensive, raina2024coaching]. Several handheld or semi-automated needle insertion devices have been developed to improve workspace, accuracy, and force resistance [slocum2014methods, cheng2019hand, ikhsan2018assistive]. For example, CathBot [cheng2019hand] enables clinicians to initiate insertion by pushing a handle, after which the device autonomously completes cannulation. While effective for peripheral vascular access on the forearm, its design has limited CVC utility due to anatomical complexities of neck, chest, or groin sites. Other robotic systems demonstrated teleoperated insertion, initiated from manually marked regions [zevallos2021toward, zevallos2025automatic] or vessel boundaries [chen2021ultrasound]. Efforts have recently shifted towards autonomy with advanced robotics, imaging, and sensor-guided needles [koskinopoulou2023robotic], reinforced by the first-in-human image-guided system in [herlihy2025centrally], showing 100% success with no complications in 19 attempts. Similarly, Hong et al. [hong2004ultrasound] developed a percutaneous insertion robot compensating for patient motion via real-time visual-servoing. Neubach et al. [neubach2009ultrasound] introduced closed-loop steering using ultrasound-estimated tip position with tissue stiffness modeling. Xu et al. [xu2010three] proposed slice-guidance interpolating preoperative scans to compensate liver motion. Chatelain et al. [chatelain20153d] used particle filtering to track flexible needles for real-time closed-loop control. While these approaches [hong2004ultrasound, neubach2009ultrasound, xu2010three, chatelain20153d] advance motion compensation, they lack integrated intra-operative planning for full autonomy.
@@
Zettinig et al. [zettinig20173d] annotated pre-interventional plans in CT/MRI and proposed a 3D visual-servoing framework with multi-modal registration to update needle targets under motion. Kojcev et al. [kojcev2016dual] developed a dual-arm ultrasound-guided system integrating force-controlled acquisition, registration, vision-based control, and tracking, though planning still relied on manual region selection. Bell et al. [bell2018photoacoustic] and Gubbi et al. [gubbi2021deep] demonstrated photoacoustic-based visual servoing. Cheng et al. [cheng2016venipuncture] used impedance sensing to detect vascular entry. Koskinopoulou et al. [koskinopoulou2023dual] combined ultrasound and bioimpedance in a dual-arm system for vessel localization and venipuncture detection. While these systems [zettinig20173d, bell2018photoacoustic, gubbi2021deep, kojcev2016dual, cheng2016venipuncture] demonstrate end-to-end procedures, scanning was typically initialized from predefined poses or required experts to review intra-operative CT/MRI images.
@@
    First robotic pipeline that autonomously performs key CVC steps: intra-operative planning, scanning, anatomical localization, and needle placement
@@
    A novel method to predict the optimal scanning region and trajectory using only a single depth image from an RGB-D camera.
@@
    A fully automated robotic scanning and vascular reconstruction for localization of vein and artery, followed by supervised needle guidance.
@@
    Feasibility demonstration of the system on a near-realistic CVC phantom for IJ vein access across 10 diverse clinical scenarios.
@@ 2.1 Experimental testbed
The experimental testbed (Fig. 1) includes: (1) a $6$-DoF UR10e robotic arm (Universal Robots, Denmark) with $10$ kg payload and inbuilt force-torque sensor; (2) a C3 HD3 curved ultrasound probe (Clarius, Canada) mounted via extended attachment (although linear probes are preferred for CVC, we used a convex probe for its wider field of view and availability during prototyping); (3) an Ubuntu PC and a laptop with NVIDIA RTX 4090 (16GB) and 1090 (12GB) for system control and deep learning-based feedback; (4) a RealSense D415 RGB-D camera (Intel, USA) for localization; (5) a custom 2-DoF needle insertion mechanism adapted from [chen2021ultrasound], providing linear and angular control, compatible with any robotic arm with adequate payload, equipped with an 18-gauge, 15 cm bevel-tipped needle (Cook Medical, USA) to constantly align its trajectory with the ultrasound plane, ensuring continuous visualization (unlike manual procedures, which rely on clinician skill to infer the needle tip); and (6) a CVC Blue Phantom (CAE, USA) for validation, offering realistic upper thorax and neck anatomy, including major vessels and landmarks for CVC. Note that the communication across system components was established via ROS2 (Robot Operating System).
Let $\boldsymbol{T}_{1}^{2}$ denote the homogeneous transformation matrix that expresses the pose of frame 1 relative to frame 2. The frames used in the system are: robot’s base (B), end-effector (E), camera (C), ultrasound probe (U), remote center-of-motion (R), and ultrasound image (I) (see Fig. 1). The 3D position in world coordinate system ($p_{B}$) of a given pixel in the 2D ultrasound image ($p_{I}$) is calculated using:
@@
where $\boldsymbol{T}_{I}^{U}$ is the transformation from the ultrasound probe frame to the pixel values in the image frame and is given by
$\boldsymbol{T}_{I}^{U}=\begin{bmatrix}S_{x}&0&0&0\\ 0&S_{y}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\begin{bmatrix}I_{x}\\ I_{y}\\ 0\\ 1\end{bmatrix}$ (2)
where $S_{x}=S_{y}=0.15436~mm$ is the pixel spacing (scaling factor) provided by the Clarius Mobile Health software development kit and $(I_{x},I_{y})$ is the given pixel value in $x$- and $y$-direction of the image frame.
@@ 2.2 Pipeline
Our pipeline is shown in Fig. 2. During initialization, system calibration was performed by estimating $\boldsymbol{T}_{C}^{B}$ via hand-eye calibration and $\boldsymbol{T}_{U}^{E}$ using the classical $AX=XB$ method [kim2013ultrasound], with a cross-wire phantom in a water bath. The region encompassing IJ vein and carotid artery was then autonomously identified from depth images, followed by robotic ultrasound scanning, vessel segmentation, 3D reconstruction, insertion point selection, and closed-loop needle guidance for puncture.
@@ 2.3 Intra-operative planning
2.3.1 Overview
Clinicians often define a triangular region using the suprasternal notch (SN), sternocleidomastoid (SCM) midpoint, and lateral clavicle (LC), known as Sedillot’s triangle, to guide IJ vein access [kosnik2024anatomical]. Inspired by this, we used SN, LC, and anterior hyoid bone tip (HT) to define a similar region, with HT chosen for its prominence and absence of SCM in the phantom. This region ensured presence of the IJ vein. These Landmarks were identified using a deep learning model, which were then used to plan the scanning path, as described below.
@@ 2.3.2 Anatomical landmark prediction network
The anatomical landmark prediction network required a depth image of the neck–torso region. Synchronized depth and color images were first captured, followed by HSV thresholding to segment the head and torso of the phantom, then depth filtering for a refined 3D point cloud. This segmented point-cloud was the network’s input (Fig. 3(a)). Following [teixeira2023automated], we trained a Dense-UNet [li2018h] to predict internal landmarks from depth images. The encoder–decoder architecture had four dense blocks of four convolutional layers (Fig. 3(b)), with batch normalization and ReLU after each convolution except the final layer.
@@
The dataset comprised 4,912 depth images of real humans with diverse habitus, sex, and skin tone. Landmarks were manually annotated using co-registered RGB-D data and verified against anthropometric measurements. The model was trained for 300 epochs with Adam optimizer (lr = 0.001, batch = 32) minimizing Adaloss, predicting 33 landmarks, including three critical to CVC, with additional landmarks to improve accuracy.
@@ 2.3.3 Scanning region and path estimation
The scanning region was defined by a triangular region formed by 3 predicted landmarks (SN, LC, and HT). This region was projected onto the phantom surface using the Open3D Python library, as shown in Fig. 3(c). To scan within this region, we defined straight-line paths from the triangle’s base (SN–LC) to its apex (HT), aligning with the IJ vein’s orientation. Initial paths connected HT to LC and HT to SN. Additional scanning paths were created by dividing the SN–LC line into equal intervals and connecting lines from HT to these points, yielding a dense scan pattern for accurate vessel detection (Fig. 3(d)). Scanning was initiated along the midpoint-to-HT path because this path aligned with the expected anatomical course of the IJ vein, offering a clinically representative initial view. This path also provided a reliable starting point across different patient positioning due to its geometric symmetry.
@@ 2.4 Robot motion planning
2.4.1 Overview
Once the scanning region was identified, the pipeline generated a robot trajectory for scanning and localizing the vein-artery, as illustrated in Fig. 4. After robotic scanning, the vein was localized and centered in the ultrasound image to ensure consistent spatial referencing. Then the robot was re-oriented for sagittal plane insertion, which offered full needle shaft visibility. A collision check was performed using the CAD model and phantom point cloud. If a collision was detected in the sagittal orientation, the robot iteratively adjusted its pivot angle in transverse view about the point of contact by $5^{\circ}$ until a collision-free pose was found. Upon confirming clearance, the robot transitioned to the sagittal plane for needle insertion.
@@ 2.4.2 Scanning trajectory
Each pose transform $\boldsymbol{P}_{*}$ was defined as $\overrightarrow{\boldsymbol{x}},\overrightarrow{\boldsymbol{y}},\overrightarrow{\boldsymbol{z}},\overrightarrow{\boldsymbol{t}}]$, where $\overrightarrow{\boldsymbol{t}}$ is the translation (or position) vector and $\overrightarrow{\boldsymbol{x}},\overrightarrow{\boldsymbol{y}},\overrightarrow{\boldsymbol{z}}$ are unit vectors for orientation. The start and end of the scanning path define $\boldsymbol{P}_{0}$ and $\boldsymbol{P}_{n}$, respectively. The normal $\overrightarrow{\boldsymbol{z}}$ of phantom was computed using Open3D and aligned with the end-effector’s normal $\overrightarrow{\boldsymbol{z}_{e}}$ using:
@@
[…]
Intermediate poses were interpolated between $\boldsymbol{P}_{0}$ and $\boldsymbol{P}_{n}$ using the UR driver. ROS2 and URScript executed the trajectory in hybrid force-position mode with $5$ N force along the $z$-axis of the probe , chosen as the minimum force required to ensure sufficient acoustic contact while minimizing tissue deformation. During scanning, a rosbag recorded poses and images pairs $(\boldsymbol{P}_{i},\boldsymbol{I}_{i})$ for each frame, described as follows to enable accurate vessel reconstruction:
@@ 2.4.3 Identifying optimal insertion plane
The acquired 2D US frames were segmented to reconstruct vessels and plan needle insertion. A deep learning model from prior work [al2025robotic] was used to extract vessel center-points and areas from each frame. Ideally, insertion should occur where both vessels (vein and artery) are clearly visible for confident and precise targeting. To determine the insertion point, the robot insertion pose ($\boldsymbol{\mathcal{P}}$) was selected from $\boldsymbol{P}_{i},\cdots,\boldsymbol{P}_{n}$ as that corresponding to the maximum combined area of the two vessels. This selection criterion is illustrated in Fig. 5. The robot was finally moved to pose $\boldsymbol{\mathcal{P}}$, and the corresponding ultrasound frame was re-segmented.
@@ 2.4.4 Vein-artery localization
Veins are difficult to distinguish from arteries, as both appear morphologically similar in ultrasound images. While clinicians typically use Doppler ultrasound to differentiate vessels based on blood flow, this method requires manual tuning, favorable imaging conditions, and expertise, making it less practical in portable or low-cost devices where Doppler functionality or detectable flow is absent. To address these limitations, our system employs a compression-based method that exploits differences in vascular compliance, whereby veins exhibit greater deformation than arteries under low external pressure.
The robot applied controlled compression by gradually increasing the contact force from $5$ N to $10$ N, then decreasing back to $5$ N, while recording an ultrasound video. Note that this compression was applied only briefly for classification. The maximum applied force of 10 N is considered clinically safe and does not induce patient discomfort. Segmented vessel areas and centroids were extracted from the recorded sequence. Then, KMeans clustering was used, which grouped vessel centroids into two clusters based on their spatial proximity. Area changes across the compression cycle were computed for each cluster, with the one showing greater deformation identified as the vein, and the other as the artery. The centroid of the vein cluster was stored as a reference. Later, at the insertion pose, the robot captured a new ultrasound frame, segmented the vessels, and identified the vein by finding the centroid closest to the saved reference, with the remaining vessel labeled as the artery.